OpenStudy (anonymous):

Which expression is (3 - 2i) (-4 + i) simplified? A) -1 - i B) 7 - 3i C) -12 + 2i D) -10 + 11i

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

do i multply?

OpenStudy (anonymous):

you simplify

OpenStudy (anonymous):

which one is the right answer do you think

OpenStudy (anonymous):

is it a)

OpenStudy (anonymous):

guess again

OpenStudy (anonymous):

c

OpenStudy (anonymous):

its d

OpenStudy (anonymous):

simplify my arse there is no such mathematical operation as "simplify" it means "multiply"

OpenStudy (anonymous):

true im wrong

OpenStudy (anonymous):

it is in fact D, do you know how to get it?

OpenStudy (anonymous):

i can show work

OpenStudy (anonymous):

-12 +3i+8i -2(i)^2 = -12 + 11i +2 = -10 + 11i

HanAkoSolo (jamierox4ev3r):

@Absoluter plz try not to provide answers. show work first :D @MyNameIsBasedGod if you didn't understand the processes, i can provide you with a recap. if you want let me know ;)

OpenStudy (anonymous):

sorry i was afk

HanAkoSolo (jamierox4ev3r):

haha thats fine but do you understand the process and stuff?

HanAkoSolo (jamierox4ev3r):

@MyNameIsBasedGod

OpenStudy (anonymous):

sorry for showing work last lol

OpenStudy (anonymous):

Yes